OpenStudy (anonymous):

Last month, Mayfield Paper Company sold 58 more than 3 times as many reams of paper as Superior Paper Products. If the two companies sold a total of sold a total of 870 reams of paper, how many reams of paper did the Mayfield Paper Company sell? Please also include work. thanks

OpenStudy (anonymous):

ash are you done?

OpenStudy (ash2326):

Let the no. of reams of paper sold by Mayfield paper company be x and no. of reams of paper sold by Superior Paper Products be y It's given that Mayfield sold 58 more then 3 times as reams of papers as Superior so \[x=3y+58\] The total sales by them is 870 \[x+y=870\] From first we have x=3y+58, substituting this in second equation so \[3y+58+y=870\] subtract both sides by 58 \[4y=870-58\] \[4y=812\] divide both sides by 4 \[y=203\] now x=3y+58 \[x=3\times 203+58\] \[x=609+58\] \[x=667\] No of reams of paepr sold by Mayfield =>x=667